the electric motor of a model train accelerates the train from rest to 0.620 m/s in 21.0cs. the total mass of the train is 875kg. find the average power delivered to the train during the acceleration.
1
Expert's answer
2019-12-27T08:56:48-0500

Work done by the electric motor


"W=\Delta E_k=\frac{mv^2}{2}"

The average power delivered to the train


"P=\frac{W}{t}=\frac{mv^2}{2t}=\frac{875\times 0.620^2}{2\times 21.0\times 10^{-2}}=801\:\rm W"